How Many Nucleosomes in One Human X Chromosome? According to the Entrez Genome Database of the National Center for Biotechnology Information, the first human X chromosome to be sequenced contained 154,913,754 nucleotide pairs. If the DNA in this chromosome is organized into nucleosomes and the average internucleosome linker DNA contains 50 nucleotide pairs, how many nucleosomes will be present in this chromosome during interphase? How many molecules of histone H3 will be present in this X chromosome?
Answer #1

During interphase chromosome will replicate to produce 2n number of chromatids hence total number of 2*154913754 = 309827508

Each nucleotide accomodate 196 bp of DNA by 146 around the nucleosome + 50 bp linker DNA = 196 bp

Number of nucleosomes = 309827508/196 = 1580753 ( 1580752.5)

Each nucleosome composed of 2 copies of H3 histone so 1580753*2 =3161506